QUESTION 25 4 points Save Answer A guidance counselor at a local high school is interested in determining what, if any, lineaPlease show how to work the problem
Answer #1

as we are instructed to use all calculations rounded off to 3 decimal places, i am using all the data values in the output by rounding off to decimal places.

the regression equation is:-

y = 1.277 + 0.17x

the predicted value of y for percentile rank (x) = 80 is:-

\hat{y} = 1.277+(0.17*80) = 14.877

t critical value for df= (n-2 )= (4137-2) = 4135 ,alpha=0.05,both tailed test be:-

t_{\frac{\alpha}{2},n-2}=1.961

the necessary data are:-

standard error of regression (s_{e}) = 0.595

\bar{x}=80.8,x^*=80,s_{x}=16.6,n=4137,\hat{y}=14.877

the 95% prediction interval is :-

1 9+tan-2 + be + 11+-+ n (x* -T2 (n-1)s

=14.877\pm 1.961*0.595*\sqrt{1+\frac{1}{4137}+\frac{(80-80.8)^2}{(4137 -1)*16.6^2}}

=\mathbf{(13.710,16.044)}

***in case of doubt, comment below. And if u liked the solution, please like.